2)

) Beginning inventory, purchases, and sales for Item HM46 are as follows: October 1 Inventory 100 units @ $19 5 Sale 80 units 1
Answer #1

Answers

  • FIFO Method
    [a] Cost of merchandise sold on Aug = (12 units x $ 25) + (7 units x $ 28) = 300 + 196 = $ 496
    [b] Inventory on Aug 31 = 35 units x $ 28 = $ 980
  • LIFO Method

[a] Cost of merchandise sold on Oct 21 = 93 units x $ 23 = $ 2139
[b] Inventory on Oct 31 = (20 units x $ 19) + (18 units x $ 23) = 380 + 414 = $ 794